OpenStudy (anonymous):

Help please i will medal and fan: Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

it is B.

OpenStudy (anonymous):

Just plug in the values of x and y and see if you get the right answer. Like for example, the order pair (2,9) Your x-value is 2. Plug in the equation y=x+7 y=2+7=9 And your original order pair does say that y-coordinate is 9. So (2,9) satisfies your equation. Repeat for other pairs as well.

OpenStudy (anonymous):

So its B like stephanie said

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

This is not a place to just give out answers. It would be better if you understand the concept. It is B, anyway.

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

I understand what you said.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I just fanned and medal you. :)